OpenStudy (experimentx):

Fundamental Theorem of Arithmetic:: ``` Every positive integer can be written as a product of primes (possibly with repetition) and any such expression is unique up to a permutation of the prime factors. (1 is the empty product, similar to 0 being the empty sum.) ``` what is the meaning of "unique up to a permutation of the prime factors" and how do you prove it?

OpenStudy (anonymous):

MATH 25 CLASS 5 NOTES, SEP 30 2011 Contents 1. Prime numbers and prime factorization 1 Quick links to denitions/theorems The Fundamental Theorem of Arithmetic 1. Prime numbers and prime factorization We've spent a good amount of time discussing divisibility, solving ax + by = c in integers, and greatest common divisors. We'll now put what we've learned to good use by studying properties of prime numbers more closely. It all starts with Euclid's Lemma, which I'll repeat again: Lemma 1 (Euclid's Lemma, 2.1b, 2.2). Let p be a prime. If pjab, then pja or pjb. More generally, if pj(a1 : : : an), then pjai for some i. Proof. We saw the proof for n = 2 last class. The proof for general n uses induc- tion. Suppose we know this statement for a given n. We want to prove the corre- sponding statement for n + 1 terms; that is, we want to prove the statement that if pj(a1 : : : an+1), then pjai for some i. Let a = a1 : : : an; b = an+1. Then pjab =) pja or pjb. If pjan+1, we are done. If pja = (a1 : : : an), use the inductive hypothesis to conclude that pjai for some 1 i n. A quick corollary of Euclid's Lemma is the following: Corollary 1 (Exercise 2.1). If p is a prime, and pjak, then pja. This is easily proven by just applying Euclid's Lemma to a1 = a; : : : ; ak = a. Why is Euclid's Lemma so important? It allows us to prove the following theorem, one of the most important in the class: Theorem 1 (The Fundamental Theorem of Arithmetic, Theorem 2.3). Let n > 1 be an integer. Then there exists a unique prime power factorization n = pe1 1 : : : pek k ; where p1; : : : ; pk are distinct prime numbers, and ei > 0 are positive integers. (When we say this factorization is unique, we really mean unique up to permutation of the prime-power factors. For instance, we consider 12 = 22 31 to be the same factoriza- tion as 12 = 31 22, since we just moved around the powers of 2; 3.)

OpenStudy (experimentx):

looks like copy and paste ... i was looking for discussion.

OpenStudy (anonymous):

there is no copy and paste that was a writing piece my brother wrote

OpenStudy (experimentx):

you mean you copied your brother's note lol

OpenStudy (anonymous):

well yup because he wrote it on my computer so that's wat happens lol

OpenStudy (experimentx):

I am aware of the theorem and proof ... there is a term "unique up to a permutation of the prime factors" I cannot decipher what it is trying to tell.

Parth (parthkohli):

It possibly means this: The prime factorization of \(N\) is unique and we consider \(N = a^x \cdot b^y \cdot c^z = b^y \cdot a^x \cdot c^z = c^z \cdot a^x \cdot b^y\), i.e., all the --permutations-- of writing the product are considered to be the same. They do not destroy the uniqueness of the prime factorization of \(N\).

ganeshie8 (ganeshie8):

prime factorization of 6 can be : 2*3 or 3*2

ganeshie8 (ganeshie8):

the prime factorization is not unique if order is considered

OpenStudy (experimentx):

why not just write "permutation of prime factorization is unique" why write the term "up to"??

Parth (parthkohli):

English is a funny language.

OpenStudy (anonymous):

tru that ^

ganeshie8 (ganeshie8):

ahh i saw that phrase "upto" in many places... let me see if i can recollect one more example of using "upto"...

OpenStudy (experimentx):

haha ... very funny

OpenStudy (experimentx):

should have posted in English section ...

Parth (parthkohli):

I think it means something along the lines of "up to considering the fact that different permutations are different things, they're not considered to be unique... but even after all that, they *are* unique."

ganeshie8 (ganeshie8):

here : http://ocw.mit.edu/courses/mathematics/18-02-multivariable-calculus-fall-2007/video-lectures/18_022007L07.pdf two different examples... search `up to` :)

OpenStudy (experimentx):

I is very very confused. Engliesh is such a wired language ... lol

Parth (parthkohli):

"any such expression is unique up to a permutation of the prime factors" This means that they're defining "unique" up to a point that different permutations of writing the prime factorization will be the same. Meaning that the expression \(a^n \times b^m\) and \(b^m \times a^n\) are considered to be the same.

OpenStudy (experimentx):

hold on a sec ... lemme decipher the paper ganishe posted.

OpenStudy (experimentx):

``` That is the solution. And any multiple of that is a solution. If you like to neatly simplify them you could say negative one, one, negative two. If you like larger numbers you can multiply that by a million. That is also a solution. Any questions about that? Yes? That is correct. If you pick these two guys instead, you will get the same solution. Well, up to a multiple. It could be if you do the cross-product of these two guys you actually get something that is a multiple ```

OpenStudy (experimentx):

is that transcript of Dennis Aurox @ganeshie8

ganeshie8 (ganeshie8):

Yes ! he uses `up to` a lot lol, atleast one time in each lecture :P

Parth (parthkohli):

lol, I lost it at "the cross-product of these two guys"

OpenStudy (experimentx):

I cannot trust that guy ... I speak better English than him lol. he is heavily french accented. JK

ganeshie8 (ganeshie8):

yes lol :) he even has a translator for composing his class lectures

OpenStudy (experimentx):

no american online??

OpenStudy (experimentx):

haha ... if he probably says then it must be correct ... after all he is professor at MIT.

ganeshie8 (ganeshie8):

i just love that guy lol, i think he was using `up to` word in the context of saying something like below : taking cross product of any two vectors lying on a same plane gives the same normal vector up to scaling...

OpenStudy (experimentx):

looks like someone posted the same question here. http://math.stackexchange.com/questions/473420/what-is-the-difference-between-being-unique-unique-up-to-isomorphism-and-unique

OpenStudy (experimentx):

lol ... these days, even you can't be the first stupid on internet. I find my problem solved even before I post. the first sentence of second paragraph solves my question. who wants the medal?

OpenStudy (experimentx):

let's make 2x2 for both of you!!

ganeshie8 (ganeshie8):

hahaha that hilarious you cant be the first stupid on internet xD give it to @ikatouni expX

Parth (parthkohli):

ikatouni = best answer ever.

OpenStudy (experimentx):

well well ... maybe one of you help him :) i hate to unto.

OpenStudy (anonymous):

exactly were is my medal